# Special Hamiltonian $S^{1}$-actions on symplectic 4-manifolds

### v3

@@ 2 Special Hamiltonian $S^{1}$-action
The unitary group $U(2)$ acts transitively on $\Lambda(\mathbb{C}^{2})$ with isotropy group $O(2)$, i.e., $\Lambda(\mathbb{C}^{2})\cong U(2)/O(2)$. By picking a unitary basis field $L$ can be represented by a loop (or half of a loop if $L$ not orientable) of unitary matrices $A$ in $U(2)$. The Maslov index $\mu(\tau)=\mu(L,\Phi)$ is then defined to be the degree of the composed map
[…]
Several remarks on $\mu(L,\Phi)$ are in order:
@@
    If $M$ is 1-connected then $\mu(L,\Phi)$ is always defined since $C$ is the boundary of some disc.
@@
    If $c_{1}(M,\omega)=0$ Then $\mu(L,\Phi)$ is independent of the class $[\psi]\in\pi_{2}(M,C)$ and hence the choice of $\psi$.
@@
Let $G$ be an $S^{1}$-group of Hamiltonian diffeomorphisms acting on $M$ effectively with nonempty fixed point set $Fix(G)$. Let $h\in C^{\infty}(M,\mathbb{R})$ denote the moment map of $G$, and $X=X_{h}$ the Hamiltonian vector field on $M$ which generates the $G$-action so that
@@
$h:M\to\mathbb{R}$ is also called a Hamiltonian function of $(M,\omega)$. For $p\not\in Fix(G)$ let $C:=Orb_{G}(p)$ denote the $G$-orbit of $p$. Take a nonzero $v\in T_{p}M/T_{p}C$ so that the plane $X|_{p}\wedge v$ spanned by $X|_{p}$ and $v$ is Lagrangian. Extend $X|_{p}\wedge v$ to an oriented $G$-invariant Lagrangian plane field $L_{p}$ along $C$ via the $G$-action. The Maslov index $\mu(L_{p})$ is independent of the choice of nonzero $v\in(T_{p}M/T_{p}C)$ with $\omega_{p}(X,v)=0$. Also $\mu(L_{p})=\mu(L_{q})$ for $q\in Orb_{G}(p)$. We say that the $G$ is special if
$\mu(L_{p})=0,\quad p\not\in Fix(G).$
@@
$\textsf{G}:=\Big{\{}\textsf{G}_{\theta}=\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}\Big{|}\theta\in\mathbb{R}/2\pi\mathbb{Z}\Big{\}}\subset SU(2)$
is a semi-free special Hamiltonian $S^{1}$-group acting on $\mathbb{C}^{2}$ with $Fix(\textsf{G})=\{0\}$.
It is easy to see that this G-action is semi-free with $Fix(\textsf{G})=\{0\in\mathbb{C}^{2}\}$, and is Hamiltonian with moment map
@@
and the corresponding Hamiltonian vector field is
@@
Maslov condition.  We claim that G is special: Since the set of nontrivial orbits of G is connected and the Maslov index is discrete, it suffices to show that the Maslov condition holds for a typical orbit of G, say the loop $C:=Orb_{\textsf{G}}(p)$ with $p=(1,0)\in\mathbb{C}^{2}$,
@@
Then $\dot{C}(0)=X_{\textsf{h}}|_{p}=\partial_{x_{2}}$. Take $v=-\partial_{x_{1}}|_{p}\in T_{p}\mathbb{C}^{2}$. Apply G to $X_{\textsf{h}}|_{p}\wedge v=\partial_{x_{2}}\wedge(-\partial_{x_{1}})|_{p}$. The resulting Lagrangian plane field along $C$ is $L:=\partial_{x_{1}}\wedge\partial_{x_{2}}$. The pair $(\partial_{x_{1}},\partial_{x_{2}})$ is a unitary basis field of the tangent bundle $T\mathbb{C}^{2}$, also a unitary basis field of $L$. Therefore under the map $\tau$ as in (1) $L$ can be represented by the constant loop $\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\subset U(2)$ with constant determinant $1$, hence the Maslov index of $L$ is $\mu(L)=0$. This confirms that the G-action on $\mathbb{C}^{2}$ is special.
Morse index of h at 00.  Note that $d\textsf{h}$ vanishes at $\{0\}=Fix(\textsf{G})$. The Hessian matrix of h at the point 00, with respect to the basis $(\partial_{x_{1}},\partial_{x_{2}},\partial_{y_{1}},\partial_{y_{2}})$, is
@@
which is nondegenerate with eigenvalues $-1,-1,1,1$, so h is a Morse function on $\mathbb{C}^{2}$, and the Morse index of h at $0\in\mathbb{C}^{2}$ is 2. The corresponding (-1)- and 1-eigenspaces are
@@
[…]
@@
[…]
@@
[…]
@@
For $m\in\mathbb{Z}\setminus\{0\}$
$\textsf{G}^{m}:=\Big{\{}\textsf{G}_{m\theta}=\begin{pmatrix}\cos m\theta&-\sin m\theta\\ \sin m\theta&\cos m\theta\end{pmatrix}\Big{|}\theta\in\mathbb{R}/(2\pi/m)\mathbb{Z}\Big{\}}\subset SU(2)$
is a special Hamiltonian $S^{1}$-group acting on $\mathbb{C}^{2}$ with weight $m$, $Fix(\textsf{G}^{m})=\{0\}$, but it is not semi-free on $\mathbb{C}^{2}$ as $stab_{p}=\mathbb{Z}_{|m|}$ if $p\neq 0$.
@@ 3 Some topological constraints imposed by Maslov condition
It turns out that at every isolated fixed point $p\in Fix(G)$ of a special Hamiltonian $S^{1}$-space $(M,\omega,G)$, the $G$-action is Hamiltonian isotopic to the standard G-action in a Darboux chart of $p$, up to a weight $m_{p}\in\mathbb{Z}\setminus\{0\}$.
@@
Let $(M,\omega,G)$ be a special Hamiltonian $S^{1}$-space with moment map $h$, and $p\in Fix(G)$ an isolated fixed point of the $G$-action.
@@
    At $p$ the $G$-action induces a linear symplectic $S^{1}$-action on the tangent space $T_{p}M$ which can be identified with $\textsf{G}\subset SU(2)$ up to a weight factor $m\in\mathbb{Z}\setminus\{0\}$, with respect to a suitable symplectic Darboux chart of $p\in M$.
@@
    There is a $G$-invariant chart $(U,x^{\prime}_{1},x^{\prime}_{2},y^{\prime}_{1},y^{\prime}_{2})$ centered at $p$ and $G$-equivariant with respect to a linear action of $G$ on $\mathbb{R}^{4}\cong\mathbb{C}^{2}$ such that $\omega|_{U}=\sum_{k=1}^{2}dx^{\prime}_{k}\wedge dy^{\prime}_{k}$.
@@
(i). For simplicity we assume that $G$ acts effectively near $p$ and with weight ±$\pm 1$. Parametrize $G$ by $t\in\mathbb{R}/2\pi\mathbb{Z}$ so that $G_{0}$ is the identity map. Fix $p\in Fix(G)$. By Darboux-Weinstein Theorem we can symplectically identify a small open neighborhood $U\subset M$ of $p$ with a open neighborhood of the origin $0\in(\mathbb{R}^{4},\omega_{0})$ of the standard symplectic $\mathbb{R}^{4}$ via a local diffeomorphism $\phi:(U,\omega)\to(\phi(U)\subset\mathbb{R}^{4},\omega_{0})$ with $\phi(p)=0$ and $\phi^{*}\omega_{0}=\omega$. Then the linear map $A_{t}:=(G_{t})_{*}|_{p}:T_{p}M\to T_{p}M$ is symplectic, i.e., $A_{t}\in Sp(4,\mathbb{R})$ for all $t$, with $A_{0}$ being the identity map. Since $G_{t}\circ G_{s}=G_{t+s}=G_{s}\circ G_{t}$ for all $s,t$, we have $A_{t}A_{s}=A_{s}A_{t}$ for all $s,t$ as a result. I.e., $A_{t}$, $t\in\mathbb{R}/2\pi\mathbb{Z}$, is an $S^{1}$-subgroup of $Sp(4,\mathbb{R})$. Up to conjugation (or equivalently a choice of $\omega_{0}$-compatible complex structure) we may assume that $A$ is an $S^{1}$-subgroup of the unitary group $U(2)=Sp(4,\mathbb{R})\cap O(4,\mathbb{R})$ with respect to the standard complex structure $J_{0}$ on $\mathbb{R}^{4}\cong\mathbb{C}^{2}$. Now $A$ is either the centralizer $S^{1}$-group $C$ of $U(2)$ or its inverse, or an $S^{1}$-subgroup of $SU(2)$.
@@
If $A=C$, then $A=\Big{\{}A_{t}=\begin{bmatrix}e^{it}&0\\ 0&e^{it}\end{bmatrix}\Big{|}t\in\mathbb{R}/2\pi\mathbb{Z}\Big{\}}$. The linear $A$-action is generated by the Hamiltonian vector field $X^{A}:=\sum_{k=1}^{2}(-y_{k}\partial_{x_{k}}+x_{k}\partial_{y_{k}})$. To check the Maslov condition, pick a point $q\not\in Fix(A)$ say, $q=(-i,0)\in\mathbb{C}^{2}$ where $X^{A}_{q}=\partial_{x_{1}}$. Let $v_{q}=\partial_{x_{2}}$. The Lagrangian plane field $L=Orb_{A}(\partial_{x_{1}}\wedge\partial_{x_{2}})$ along $\gamma:=Orb_{A}(q)$ is $L(t)=(\cos t\partial_{x_{1}}+\sin t\partial_{y_{1}})\wedge(\cos t\partial_{x_{2}}+\sin t\partial_{y_{2}})$, $t\in\mathbb{R}/2\pi\mathbb{Z}$. With its unitary basis field $(\cos t\partial_{x_{1}}+\sin t\partial_{y_{1}},\cos t\partial_{x_{2}}+\sin t\partial_{y_{2}})$ $L$ can be represented by the loop of unitary matrices $\begin{bmatrix}e^{it}&0\\ 0&e^{it}\end{bmatrix}$. By (1) the Maslov index of $L$ is $\mu(L)=4\neq 0$. So $A\neq C$. Similarly $A\neq C^{-1}$ as otherwise we would have $\mu(Orb_{A})(\partial_{x_{1}}\wedge\partial_{x_{2}})=-4\neq 0$. So $A$ is an $S^{1}$-subgroup of $SU(2)$. Since all $S^{1}$-subgroups of $SU(2)$ are conjugate we have $A=\textsf{G}$ up to a change of the complex coordinates of $\mathbb{C}^{2}$ via a special unitary linear map given by some element of $SU(2)$.
@@
As the set $Fix(G)$ is discrete without accumulation points we may choose for each $p\in Fix(G)$ a symplectic Darboux chart $U_{p}$ such that $U_{p}\cap U_{q}=\emptyset$ for $p,q\in Fix(G)$ with $p\neq q$. The above result about $A$ applies to every $p\in Fix(G)$ as a result. This finishes the proof of (i).

### v4

@@ 2 Special Hamiltonian $S^{1}$-action
The unitary group $U(2)$ acts transitively on $\Lambda(\mathbb{C}^{2})$ with isotropy group $O(2)$, i.e., $\Lambda(\mathbb{C}^{2})$ is conjugate to $U(2)/O(2)$. By picking a unitary basis field $L$ can be represented by a loop (or half of a loop if $L$ not orientable) of unitary matrices $A$ in $U(2)$. The Maslov index $\mu(\tau)=\mu(L,\psi)$ is then defined to be the degree of the composed map
(1) ${\det}^{2}\circ\tau:S^{1}\overset{\tau}{\to}U(2)\overset{\det^{2}}{\to}U(1)\cong S^{1}.$
Several remarks on $\mu(L,\psi)$ are in order:
@@
    If $M$ is 1-connected then $\mu(L,\psi)$ is always defined since $C$ is the boundary of some disc.
@@
    If $c_{1}(M,\omega)=0$ Then $\mu(L,\psi)$ is independent of the class $\psi]\in\pi_{2}(M,C)$ and hence the choice of $\psi$.
@@
Let $\Phi$ be an $S^{1}$-group of Hamiltonian diffeomorphisms acting on $M$ effectively with nonempty fixed point set $Fix(\Phi)$. Let $h\in C^{\infty}(M,\mathbb{R})$ denote the moment map of $\Phi$, and $X=X_{h}$ the Hamiltonian vector field on $M$ which generates the $\Phi$-action so that
@@
$h:M\to\mathbb{R}$ is also called a Hamiltonian function of $(M,\omega)$. For $p\not\in Fix(\Phi)$ let $C:=Orb_{\Phi}(p)$ denote the $\Phi$-orbit of $p$. Take a nonzero $v\in T_{p}M/T_{p}C$ so that the plane $X|_{p}\wedge v$ spanned by $X|_{p}$ and $v$ is Lagrangian. Extend $X|_{p}\wedge v$ to an oriented $\Phi$-invariant Lagrangian plane field $L_{p}$ along $C$ via the $G$-action. The Maslov index $\mu(L_{p})$ is independent of the choice of nonzero $v\in(T_{p}M/T_{p}C)$ with $\omega_{p}(X,v)=0$. Also $\mu(L_{p})=\mu(L_{q})$ for $q\in Orb_{\Phi}(p)$. We say that the $\Phi$ is special if
$\mu(L_{p})=0,\quad p\not\in Fix(\Phi).$
@@
$\Theta:=\Big\{\Theta_{\theta}=\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}\Big|\theta\in\mathbb{R}/2\pi\mathbb{Z}\Big\}\subset SU(2)$
is a semi-free special Hamiltonian $S^{1}$-group acting on $\mathbb{C}^{2}$ with $Fix(\Theta)=\{0\}$.
It is easy to see that this $\Theta$-action is semi-free with $Fix(\Theta)=\{0\in\mathbb{C}^{2}\}$, and is Hamiltonian with moment map
@@
the corresponding Hamiltonian vector field is
@@
Maslov condition. We claim that $\Theta$ is special: Since the set of nontrivial orbits of $\Theta$ is connected and the Maslov index is discrete, it suffices to show that the Maslov condition holds for a typical orbit of $\Theta$, say the loop $C:=Orb_{\Theta}(p)$ with $p=(1,0)\in\mathbb{C}^{2}$,
@@
Then $\dot{C}(0)=X_{\textsf{h}}|_{p}=\partial_{x_{2}}$. Take $v=-\partial_{x_{1}}|_{p}\in T_{p}\mathbb{C}^{2}$. Apply $\Theta$ to $X_{\textsf{h}}|_{p}\wedge v=\partial_{x_{2}}\wedge(-\partial_{x_{1}})|_{p}$. The resulting Lagrangian plane field along $C$ is $L:=\partial_{x_{1}}\wedge\partial_{x_{2}}$. The pair $(\partial_{x_{1}},\partial_{x_{2}})$ is a unitary basis field of the tangent bundle $T\mathbb{C}^{2}$, also a unitary basis field of $L$. Therefore under the map $\tau$ as in (1) $L$ can be represented by the constant loop $\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\subset U(2)$ with constant determinant $1$, hence the Maslov index of $L$ is $\mu(L)=0$. This confirms that the $\Theta$-action on $\mathbb{C}^{2}$ is special. $\Box$
Morse index of h at 0. Note that $d\textsf{h}$ vanishes at $\{0\}=Fix(\Theta)$. The Hessian matrix of h at the point 0, with respect to the basis $(\partial_{x_{1}},\partial_{x_{2}},\partial_{y_{1}},\partial_{y_{2}})$, is
@@
which is nondegenerate with eigenvalues 1,1,1,1-1,-1,$1,$ so h is a Morse function on $\mathbb{C}^{2}$, and the Morse index of h at $0\in\mathbb{C}^{2}$ is 2. The corresponding (-1)- and 1-eigenspaces are
@@
Stable/unstable submanifolds of h. Let $\nabla\textsf{h}$ denote the gradient vector field of h relative to the standard Euclidean metric on $\mathbb{R}^{4}\cong\mathbb{C}^{2}$, which is both $\omega_{0}$-compatible and G-invariant. Then
@@
Let $W_{\textsf{h}}^{s}(0)$ and $W_{\textsf{h}}^{u}(0)$ denote the stable and unstable submanifolds of of the flow of $\nabla\textsf{h}$ at $0\in\mathbb{R}^{4}\cong\mathbb{C}^{2}$ respectively. We claim that
@@
Consider the point $(1,i)\in\mathbb{C}^{2}$ on $E_{-1}$, the integral curve of $\nabla\textsf{h}$ passing through the point $(1,i)$ at time $t=0$ is $\ell_{(1,i)}(t)=(e^{-t},ie^{-t})\subset E_{-1}$, the half $\mathbb{R}$-space spanned by $\nabla\textsf{h}(1,i)=-(\partial_{x_{1}}+\partial_{y_{2}})$. The $\Theta$-action preserves the space of integral curves of $\nabla\textsf{h}$. In particular $\Theta_{\pi/2}(\ell_{(1,i)})=\ell_{(-i,1)}\subset E_{-1}$ is the integral curve of $\nabla\textsf{h}$ passing through the point (i,1)(-i,1) at time $t=0$, also a half $\mathbb{R}$-space spanned by $\nabla\textsf{h}(-i,1)=\partial_{y_{1}}-\partial_{x_{2}}\subset E_{-1}$. Moreover, $Orb_{\Theta}(\ell_{(1,i)})=E_{-1}\setminus\{0\}$. So $W_{\textsf{h}}^{s}(0)=E_{-1}=\{z_{1}+iz_{2}=0\}$.
The case for $W_{\textsf{h}}^{u}(0)$ can be verified in a similar way. Consider the point $(1,-i)\in\mathbb{C}^{2}$ on $E_{1}$, the integral curve of $\nabla\textsf{h}$ passing through (1,i)(1,-i) at time $t=0$ is $\ell_{(1,-i)}(t)=(e^{t},-ie^{t})\subset E_{1}$, the half $\mathbb{R}$-space spanned by $\nabla\textsf{h}(1,-i)=\partial_{x_{1}}-\partial_{y_{2}}$. $\Theta_{\pi/2}(\ell_{(1,-i)})=\ell_{(i,1)}\subset E_{1}$ is the integral curve of $\nabla\textsf{h}$ passing through the point $(i,1)$ at time $t=0$, also a half $\mathbb{R}$-space spanned by $\nabla\textsf{h}(i,1)=\partial_{y_{1}}+\partial_{x_{2}}\subset E_{1}$. Moreover, $Orb_{\Theta}(\ell_{(1,-i)})=E_{1}\setminus\{0\}$. So $W_{\textsf{h}}^{u(0)}=E_{1}=\{z_{1}-iz_{2}=0\}$. $\Box$
@@
Observe that $\Theta^{-1}$ is the inverse action of $\Theta$ on $\mathbb{C}^{2}$, with the pair $(-X,-\textsf{h})$ as its corresponding Hamiltonian vector field and moment map instead, and orbits of the $\Theta^{-1}$-action are precisely the orbits of $\Theta$-action but with the opposite orientation. Nevertheless, $\Theta$ and $\Theta^{-1}$ act on $\mathbb{C}^{2}$ symplectically equivariant. For example, consider the symplectic diffeomorphism $f_{0}:\mathbb{C}^{2}\to\mathbb{C}^{2}$ given by
$f_{0}(x_{1},y_{1},x_{2},y_{2})=(-x_{1},-y_{1},x_{2},y_{2}),$
$f_{0}$ is a linear symplectic map which can be represented by the unitary matrix
$\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}$,
and $f_{0}=f_{0}^{-1}$. A direct computation shows that
$f_{0}^{-1}\circ\Theta^{-1}\circ f_{0}=\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}\begin{bmatrix}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{bmatrix}\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}=\Theta.$
So $\Theta$ and $\Theta^{-1}$ are equivariant SHam-1 actions on $\mathbb{C}^{2}$. $\Box$
@@ 3 Some topological constraints imposed by Maslov condition
It turns out that at every isolated fixed point $p\in Fix(\Phi)$ of a special Hamiltonian $S^{1}$-space $(M,\omega,\Phi)$, the $\Phi$-action is Hamiltonian isotopic to the standard $\Theta$-action in a Darboux chart of $p$.
@@
Let $(M,\omega,\Phi)$ be a SHam1-space with moment map $h$, and $p\in Fix(\Phi)$ an isolated fixed point of the $\Phi$-action.
@@
    At $p$ the $\Phi$-action induces a linear symplectic $S^{1}$-action on the tangent space $T_{p}M$ which can be identified with $\Theta\subset SU(2)$ with respect to a suitable symplectic Darboux chart of $p\in M$.
@@
    There is a $\Phi$-invariant chart $(V,x_{1},x_{2},y_{1},y_{2})$ centered at $p$ and $\Phi$-equivariant with respect to a linear action of $\Phi$ on $\mathbb{R}^{4}\cong\mathbb{C}^{2}$ such that $\omega|_{V}=\sum_{k=1}^{2}dx_{k}\wedge dy_{k}$.
@@
(i). Parametrize $\Phi$ by $t\in\mathbb{R}/2\pi\mathbb{Z}$ so that $\Phi_{0}$ is the identity map. Fix $p\in Fix(\Phi)$. By Darboux-Weinstein Theorem we can symplectically identify a small open neighborhood $V\subset M$ of $p$ with a open neighborhood of the origin $0\in(\mathbb{R}^{4},\omega_{0})$ of the standard symplectic $\mathbb{R}^{4}$ via a local diffeomorphism $\phi:(V,\omega)\to(\phi(V)\subset\mathbb{R}^{4},\omega_{0})$ with $\phi(p)=0$ and $\phi^{*}\omega_{0}=\omega$. Then the linear map $A_{t}:=(\Phi_{t})_{*}|_{p}:T_{p}M\to T_{p}M$ is symplectic, i.e., $A_{t}\in Sp(4,\mathbb{R})$ for all $t$, with $A_{0}$ being the identity map. Since $\Phi_{t}\circ\Phi_{s}=\Phi_{t+s}=\Phi_{s}\circ\Phi_{t}$ for all $s,t$, we have $A_{t}A_{s}=A_{s}A_{t}$ for all $s,t$ as a result. I.e., $A_{t}$, $t\in\mathbb{R}/2\pi\mathbb{Z}$, is an $S^{1}$-subgroup of $Sp(4,\mathbb{R})$. Up to conjugation (or equivalently a choice of $\omega_{0}$-compatible complex structure) we may assume that $A$ is an $S^{1}$-subgroup of the unitary group $U(2)=Sp(4,\mathbb{R})\cap O(4,\mathbb{R})$ with respect to the standard complex structure $J_{0}$ on $\mathbb{R}^{4}\cong\mathbb{C}^{2}$. Now $A$ is either the centralizer $S^{1}$-group $C$ of $U(2)$ or an $S^{1}$-subgroup of $SU(2)$.
@@
If $A=C$ then $A=\Big\{A_{t}=\begin{bmatrix}e^{it}&0\\ 0&e^{it}\end{bmatrix}\Big|t\in\mathbb{R}/2\pi\mathbb{Z}\Big\}$. The linear $A$-action is generated by the Hamiltonian vector field $X^{A}:=\sum_{k=1}^{2}(-y_{k}\partial_{x_{k}}+x_{k}\partial_{y_{k}})$. To check the Maslov condition, pick a point $q\not\in Fix(A)$ say, $q=(-i,0)\in\mathbb{C}^{2}$ where $X^{A}_{q}=\partial_{x_{1}}$. Let $v_{q}=\partial_{x_{2}}$. The Lagrangian plane field $L=Orb_{A}(\pm\partial_{x_{1}}\wedge\pm\partial_{x_{2}})$ along $\gamma:=Orb_{A}(q)$ is $L(t)=(\cos t\partial_{x_{1}}+\sin t\partial_{y_{1}})\wedge(\cos t\partial_{x_{2}}+\sin t\partial_{y_{2}})$, $t\in\mathbb{R}/2\pi\mathbb{Z}$. With its unitary basis field $(\cos t\partial_{x_{1}}+\sin t\partial_{y_{1}},\cos t\partial_{x_{2}}+\sin t\partial_{y_{2}})$ $L$ can be represented by the loop of unitary matrices $\begin{bmatrix}e^{it}&0\\ 0&e^{it}\end{bmatrix}$. By (1) the Maslov index of $L$ is $\mu(L)=4\neq 0$. So $A\neq C$. Similarly $A\neq C^{-1}$ as otherwise we would have $\mu(Orb_{A})(\partial_{x_{1}}\wedge\partial_{x_{2}})=-4\neq 0$. So $A$ is an $S^{1}$-subgroup of $SU(2)$. Since all $S^{1}$-subgroups of $SU(2)$ are conjugate we have $A=\Theta$ or $\Theta^{-1}$ up to a change of the complex coordinates of $\mathbb{C}^{2}$ via a special unitary linear map given by some element of $SU(2)$.
@@
As the set $Fix(\Phi)$ is discrete without accumulation points we may choose for each $p\in Fix(\Phi)$ a symplectic Darboux chart $U_{p}$ such that $U_{p}\cap U_{q}=\emptyset$ for $p,q\in Fix(\Phi)$ with $p\neq q$. The above result about $A$ applies to every $p\in Fix(\Phi)$ as a result. This finishes the proof of (i).